OpenStudy (anonymous):

59 ÷ 54 *56 *53 *55 *54

OpenStudy (mathstudent55):

Those are exponents, right? \(\dfrac{5^9}{5^4} \) Use the rule \(\dfrac{a^m}{a^n} = a^{m - n} \)

OpenStudy (anonymous):

yeah 5 with an exponent of 6 @mathstudent55 and the second number so on is an exponent

OpenStudy (mathstudent55):

When you divide two powers with the same base, give the same base and subtract the exponents.